OpenStudy (anonymous):

The difference of a number squared and the number is one less than four times the number.

OpenStudy (anonymous):

wait are you looking for an equation?@thecollinjc1

OpenStudy (anonymous):

x^2 - x = 4x - 1 ?

OpenStudy (anonymous):

So The Equation Would Be x^2 - x = 4x - 1?

OpenStudy (anonymous):

x^2 - x = 4x - 1